OpenStudy (anonymous):

What is the value of x? -6x=42 What is the value of x? -2x+8=-16 What is the value of x? 3(x-2)+5x=18

OpenStudy (anonymous):

1. -7

OpenStudy (anonymous):

@mergas12345 can u help me with these and are u shure ur givin me the right anwers?

OpenStudy (anonymous):

x = -7

OpenStudy (gorv):

1 divide both side by 6

OpenStudy (anonymous):

yes that's write

OpenStudy (anonymous):

-7 is the first one,GUARANTEED!!

OpenStudy (anonymous):

1 is -6 actually

OpenStudy (gorv):

2 subtacr 8 from both side and divide by -2

OpenStudy (anonymous):

2. x=12

OpenStudy (gorv):

3 open the bracket 3x-6+5x=18 8x=12 x=12/8=3/2

OpenStudy (anonymous):

calculators dont lie i just did it in my head and got -6 then i checked my work on my calculator and i got -6

OpenStudy (anonymous):

for the first one

OpenStudy (anonymous):

and 3 is 3

OpenStudy (anonymous):

so, the answer to 1 is -7, the answer to 2 is -12, and the answer to 3 is 3

OpenStudy (anonymous):

so is the first one -6 or -7

OpenStudy (anonymous):

@igivemedalsandfollow!

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

i hope i helped